The following table describes the top 6 causes of death among Americans ages 15 to 24 in 2015. There were over 44 million (44,000,000) people between the ages of 15 to 24 in the US in 2015.

Rank

Cause of Death

Number of Deaths

1

Unintentional Injury

12,514

2

Suicide

5,491

3

Homicide

4,733

4

Malignant neoplasms

1,469

5

Heart Disease

997

6

Congenital Anomalies (Birth Defects)

386

Q. What was the proportional mortality for malignant neoplasms among the top 6 causes of death? Provide a one sentence interpretation.
Answer #1

Total deaths = 12514 + 5491 + 4733 + 1469 + 997 + 386 = 25590

Proportional mortality for malignant neoplasms among the top deaths here is computed as:

= Number of deaths by malignant neoplasms / Total deaths

= 1469/25590

= 0.0574

Therefore 0.0574 is the required proportional mortality here.